# Safety through feedback in Constrained RL

### v1

@@ Appendix A Proofs
$L^{sur}-L^{mle}$ $=-\frac{1}{|P|}\sum_{\tau_{i:j}\sim P}(1-y^{safe})\left[\sum_{t=i}^{j}\log{(1-p^{safe}(s_{t},a_{t}))}-\log{(1-\prod_{t=i}^{j}p^{safe}(s_{t},a_{t}))}\right]$
$=\frac{1}{|P|}\sum_{\tau_{i:j}\sim P}(1-y^{safe})\left[\log{\left(\frac{1-\prod_{t=i}^{j}p^{safe}(s_{t},a_{t})}{\prod_{t=i}^{j}1-p^{safe}(s_{t},a_{t})}\right)}\right]$
@@ B.1 Benchmark environments:
simulator [28] based environments extend the locomotion tasks to include safety constraints. The environments can be split into two types: position based constraints and velocity based, a detailed description is provide below:
@@
    Position Based: These environments are recent benchmarks introduced in [20]. Hence, we explain them in more detail.
@@
      Blocked Swimmer: The agent controls a robot that consists of three segments joined at two points. The goal is to move the agent towards the right by as fast as possible by the application of torque via rotors connected to the joints. The reward received after each time step is proportional to its motion in the $X$-direction during that timestep. The episode ends after $1000$ timesteps is reached. It is easier for the agent to move to the right, hence to make the task challenging a cost is incurred if the $X$ coordinate of the agent $\geq 0.5$, constraining the agent to the region $\infty\leq X<0.5$.
@@
      Biased Pendulum: The agents goal is to balance a pole on a cart. Each episode ends if the pole is dropped or a maximum of $1000$ timesteps is reached. The agent receives a reward $1$ when its $X$-coordinate lies in the region $[\infty,\-0.01]$ and monotionically decreases to $0.1$ as at its $X$ coordinate increases to 00 after which it receives a constant reward of $0.1$, hence the agent is biased to move in the leftward direction. Consequently, a cost of $1$ is incurred in the region $(-\infty,-0.015]$ constraining the agent from moving too far left.
@@
    Velocity Based: These environments are a part of the Safety Gymnasium benchmarks. They add a constraint on the velocity of the Half-Cheetah, Hopper and Walker-2d agents respectively. The agent incurs a cost of $1$ whenever it exceeds a threshold of $v_{max}/2$, where $v_{max}$ is the maximum velocity that agent receives when trained by $PPO$ for $1e^{7}$ steps.
@@
[15] has emerged as a prominent benchmark when evaluating constrained RL algorithms, thus providing a challenging benchmark for cost inference. These environments consist of multiple agents such Point, Car and Doggo, listed in increasing order of the difficulty in controlling the agent. The goal is to complete a certain task using any one of these agents. In the Circle task, the agent must go around the boundary of a circle centered around the origin. However, the agent must do so while staying within a safe region which is characterized by two vertical boundaries located at $x=\pm 0.7$. A cost of $1$ is incurred every time the agent is outside the confines of this boundary ($|x|\geq 0.7$). In the Goal task, the agent must reach a designated goal area as soon as possible while simultaneously avoiding static and dynamic obstacles. A cost of $1$ is incurred every time the agent collides with these obstacles. The Push task is similar to the Goal task with the added challenge that the agent must now push a block to the goal area while simultaneously avoiding collision with obstacles present in the environment.
@@ B.2 Safe Driver
Driving based environments are emerging as benchmarks for the problem of constraint inference [20], owing to the nuanced aspect of safety when it comes to driving. We test our algorithm on the Driver simulator introduced in [19, 22] to learn safe behaviours from feedback. This environment consists of two scenarios that a self driving agent is likely to encounter on the Highway: a blocked road and a lane change. We add another scenario that represents overtaking in a two lane highway, where the agent must overtake a slower car in its lane by using the second lane. This can be dangerous as the traffic in this lane is in the other direction and hence the agent must do so in a safe manner. We provide a brief description of the environment below.
@@
The Driver environment used point-mass dynamic and consists of continuous state and action spaces. The state $s_{i}$ of a vehicle in the simulation is described by the tuple $(x,y,\phi,v)$, that consists of the agent’s position $(x,y)$, heading $\phi$ and velocity $v$. The observation of the ego vehicle consists of the concatenation of the states of all the vehicles in the scene. The environment takes two actions $(a_{1},a_{2})$ where $a_{1}$ represents the steering input $a_{2}$ represents the applied acceleration.
@@
The task is to control the ego vehicle (white in Figure 3) to reach the top of the road as soon as possible while adhering to costs on the velocity and distance to other vehicles. The trajectory of the other vehicles is fixed bar some random noise applied to its steering and acceleration inputs. The agent receives a reward $r_{t}=10(y_{t}-y_{t-1})$ every timestep, incentivizing it to cut through the traffic as fast as possible. Additionally it receives a penalty $-1$ if it goes off the road, drives backward or does not stay in the middle of the lane. It also incurs a penalty of $-100$ for for a collision. The episode is terminated if the agent reaches the top of the road or collides with another vehicle. We show that although such constraints can be incorporated into the reward function, it still leads to highly unsafe behaviour, highlighting the need for incorporating an additional cost function. The PPO agent is trained on this reward function.
@@
In the CMDP settings, the cost function absorbs some of the constraints from the reward, i.e, the agent receives a cost of $1$ when it goes off the road or drives backward. Additional costs are incurred when the agent crosses a speed limit $v_{max}$ or gets too close to any nearby vehicle. The ego car is deemed to close to a neighbouring vehicle if $\exp{-b(c_{1}d_{x}^{2}+c_{2}d_{y}^{2})+ba}\geq 0.4$, where $a=0.01$, $b=30$, $c_{1}=10$, $c_{2}=2$ and $d_{x}$ and $d_{y}$ represent the distance of the nearby vehicle from the ego vehicle.
Appendix C Experiments
C.1 Hyperparameters
@@
We conducted the experiments on a cluster containing $4$ NVIDIA RTX A5000 GPUs and $96$ core CPUs. The experiments took an approximate of two weeks to run. The detailed hyperparameters used in the experiments are reported in Table 3.
@@ C.2 Main Experiments
Table 4: Number of Trajectories shown to the Evaluator for Feedback
@@
C.3 Novelty based Sampling
Figure 6 shows the implicit decreasing schedule encountered when using novelty based sampling in various environments. Figure 7 shows additional comparison with other querying mechanism in the Biased Pendulum environment.
@@ C.3 Novelty based Sampling
C.4 Number of Queries generated
@@
C.5 Complexity of Constraint Inference
@@
Figure 8 illustrates the challenge of inferring constraints when cost violations are sparse. Due to the brief interaction periods with obstacles in the environment, learning from feedback over longer horizons makes the problem of assigning credit to the unsafe states challenging. Therefore, feedback must be elicited at the state level becomes in such scenarios.

### v2

@@ Appendix A Proofs
$L^{sur}-L^{mle}$ $=-E_{(\tau_{i:j},y^{safe})\sim P}(1-y^{safe})\Bigg{[}\sum_{t=i}^{j}\log{(1-p^{safe}(s_{t},a_{t}))}$
$\quad-\log{(1-\prod_{t=i}^{j}p^{safe}(s_{t},a_{t}))}\Bigg{]}$
$=E_{(\tau_{i:j},y^{safe})\sim P}(1-y^{safe})\left[\log{\left(\frac{1-\prod_{t=i}^{j}p^{safe}(s_{t},a_{t})}{\prod_{t=i}^{j}1-p^{safe}(s_{t},a_{t})}\right)}\right]$
@@ B.1 Benchmark environments:
[30]-based environments extend traditional locomotion tasks by incorporating safety constraints. These environments are categorized into two types: position-based and velocity-based constraints. A detailed description of each is provided below:
@@
    Position Based: These environments, recently introduced as benchmarks in [22], are explained in more detail below.
@@
      Blocked Swimmer: The agent controls a robot composed of three segments joined at two points. The objective is to move the robot to the right as quickly as possible by applying torque through rotors at the joints. The agent receives a reward at each time step, proportional to its displacement in the $X$-direction. The episode concludes after 1000 time steps. To increase the challenge, a cost is imposed when the agent’s $X$-coordinate exceeds 0.5, constraining it to the region $-\infty\leq X<0.5$.
@@
      Biased Pendulum: The agent’s goal is to balance a pole on a cart. Each episode ends if the pole falls or after a maximum of 1000 time steps. The agent receives a reward of 1 when its $X$-coordinate is within the range $(-\infty,-0.01]$, which monotonically decreases to 0.1 as the $X$-coordinate approaches 0. Beyond this, the agent receives a constant reward of 0.1, encouraging movement in the leftward direction. To prevent excessive leftward movement, a cost of 1 is incurred when the $X$-coordinate is in the region $(-\infty,-0.015]$, constraining the agent from moving too far left.
@@
    Velocity Based: These environments are part of the Safety Gymnasium benchmarks, introducing a velocity constraint for the Half-Cheetah, Hopper, and Walker-2d agents. The agent incurs a cost of 1 whenever it exceeds a velocity threshold of $v_{max}/2$, where $v_{max}$ is the maximum velocity achieved by the agent when trained with PPO for $1\times 10^{7}$ steps.
@@
[17] has emerged as a prominent benchmark for evaluating constrained reinforcement learning algorithms, providing a challenging framework for cost inference. The environments feature multiple agents, including the Point, Car, and Doggo, listed in increasing order of difficulty in controlling the agent. The objective is to complete specific tasks using one of these agents.
In the Circle task, the agent must navigate around the boundary of a circle centered at the origin while remaining within a safe region defined by two vertical boundaries at $x=\pm 0.7$. A cost of 1 is incurred each time the agent exceeds this boundary ($|x|\geq 0.7$).
In the Goal task, the agent must reach a designated goal area as quickly as possible while avoiding both static and dynamic obstacles, incurring a cost of 1 for each collision.
The Push task builds upon the Goal task by adding the challenge of pushing a block to the goal area while still avoiding collisions with obstacles in the environment.
@@ B.2 Safe Driver
Driving-based environments are emerging as key benchmarks for addressing the problem of constraint inference [22], particularly due to the complex safety considerations involved in driving. We evaluate our algorithm using the Driver simulator introduced in [21, 24] to learn safe behaviors from feedback. This environment encompasses two scenarios that a self-driving agent is likely to encounter on the highway: a blocked road and a lane change. Additionally, we introduce a third scenario that involves overtaking on a two-lane highway, where the agent must safely pass a slower car in its lane by utilizing the second lane. This maneuver poses risks, as traffic in this lane is moving in the opposite direction, necessitating careful execution. Below, we provide a brief description of the environment.
@@
The Driver environment utilizes point-mass dynamics and features continuous state and action spaces. The state of a vehicle in the simulation is represented by the tuple $(x,y,\phi,v)$, which includes the agent’s position $(x,y)$, heading $\phi$ and velocity $v$. The observation of the ego vehicle is formed by concatenating the states of all vehicles in the scene. The environment allows for two actions $(a_{1},a_{2})$ where $a_{1}$ represents the steering input $a_{2}$ denotes the applied acceleration.
@@
The task is to control the ego vehicle (white in Figure 5) to reach the top of the road as soon as possible while adhering to costs on the velocity and proximity to other vehicles. The trajectories of the other vehicles are fixed, with some random noise applied to their steering and acceleration inputs. The agent receives a reward $r_{t}=10(y_{t}-y_{t-1})$ at each time step, incentivizing it to navigate through traffic rapidly. Additionally, it incurs a penalty of $-1$ for going off the road, driving backward, or failing to stay in the center of the lane. A collision results in a penalty of $-100$. The episode terminates when the agent reaches the top of the road or collides with another vehicle. Although these constraints can be integrated into the reward function, they can still lead to unsafe behaviors, highlighting the necessity for an additional cost function. The PPO agent is trained using this reward function.
@@
In the CMDP settings, the cost function incorporates some of the constraints from the reward function. For instance, the agent incurs a cost of $1$ when it goes off the road or drives backward. Additional costs are incurred when the agent crosses the speed limit $v_{max}$ or gets too close to any nearby vehicles. The ego car is deemed to close to a neighbouring vehicle if $\exp{-b(c_{1}d_{x}^{2}+c_{2}d_{y}^{2})+ba}\geq 0.4$, where $a=0.01$, $b=30$, $c_{1}=10$, $c_{2}=2$ and $d_{x}$ and $d_{y}$ represent the distance of the nearby vehicle from the ego vehicle.
Appendix C Related Work
Constrained RL
The Constrained Markov Decision Process (CMDP) framework [4] has emerged as a valuable tool in safety-critical settings, offering a clear separation between task-related information (rewards) and safety-related information (costs). This separation simplifies environment specification and allows for easier transfer of cost functions across similar environments. While previous research has extensively addressed this problem [1, 35, 34, 16], these studies typically assume that the cost function is known. However, this assumption can be limiting when designing the cost function is complex or costly to evaluate. Consequently, there is a need to infer the cost function from external data in such cases.
Cost Inference
Historically, cost inference has primarily depended on two types of data: (1) evaluator feedback and (2) constraint abiding demonstrations obtained from an expert.
Previous work on learning from feedback often relies on limiting assumptions such as deterministic transitions and smoothness [28, 32, 11] or linearity assumptions [5] on the cost function. Moreover, none of these approaches address the problem of actively querying humans to minimize the required feedback. Although this has been theoretically explored in [8], the proposed algorithm has not yet been adapted for complex environments. Additionally, these methods are typically limited to feedback on individual state-action pairs, which can be expensive when relying on human evaluators. Our RLSF approach addresses these limitations by reducing the amount of feedback required and avoiding assumptions about the cost function.
Inferring cost functions from constraint-satisfying expert demonstrations has gained significant attention recently [6, 33, 22]. However, acquiring such expert demonstrations can be challenging in many scenarios. An alternative form of feedback is human interventions. For instance, [23] rely on the human evaluator to terminate the episode when the agent is about to engage in unsafe behavior, while [25] depend on the human to take control of the agent in such situations and guide it to a safe state.
Novelty-Based Sampling
Novelty estimation has been widely explored in the context of active exploration in reinforcement learning. The connection between prediction accuracy and state novelty has been extensively studied, as highlighted by [9]. Additionally, state density estimation has been investigated in works such as [29, 7, 13], where infrequently visited states are assigned an exploration bonus. Notably, [29] also uses Simhash [10] to maintain state density estimates.
Reward Inference
This work also connects to the analogous problem of reward inference in standard MDPs, which similarly rely on human feedback [12, 18, 20] and expert demonstrations [15, 14].
Appendix D Experiments
D.1 Hyperparameters
@@
We conducted the experiments on a cluster quipped with $4$ NVIDIA RTX A5000 GPUs and $96$ core CPUs. The experiments took an approximate of two weeks to run. The detailed hyperparameters utilized in the experiments are presented in Table 3.
@@ D.2 Main Experiments
Table 4: Number of trajectories presented to the evaluator for feedback
@@
D.3 Novelty based Sampling
Figure 6 shows the implicit decreasing schedule encountered when using novelty based sampling in various environments. Figure 9 shows additional comparison with other querying mechanisms in the Biased Pendulum environment. To further assess the effectiveness of novelty based sampling, we categorized the trajectories collected during a data collection round based on novelty. We then evaluated the model’s accuracy on the pseudo-prediction task for these two trajectory types, as shown in Figure 10. This analysis indicates that the model’s accuracy is lower for novel trajectories, corroborating the claims made in Section 4.3.
@@ D.3 Novelty based Sampling
Figure 10: Model prediction accuracy (mean $\pm$ standard error) in the surrogate task (averaged over 3 seeds) with trajectory level feedback. Predictions are made on the next $50$k steps following $250$k training steps.
D.4 Number of Queries generated
@@
D.5 Complexity of Constraint Inference
@@
Figure 11 highlights the difficulty of inferring constraints when cost violations are sparse. The brief interaction periods with obstacles in the environment make it challenging to assign credit to unsafe states when feedback is collected over long horizons. Consequently, feedback was collected at the state level in such settings. This is illustrated in Figure 12, where the overestimation bias is considerably higher in the Point Goal environment compared to Point Circle, resulting in a significant drop in performance.
D.6 Ablation on the size of the feedback buffer
We conducted an ablation study on the size of the feedback buffer, the results of which are presented in Figure 13. We can observe that the policy becomes overly conservative when the size of the feedback buffer is small (100k, which corresponds to approximately two data collection rounds). We hypothesize that this is because of two interacting factors: (a) discarding past feedback can result in forgetting, which in turn leads to inaccuracies in estimating unsafe regions, and (b) there exists a feedback loop in which the inferred cost function influences on-policy data collection. If the inferred cost function overestimates costs in certain regions, the policy will avoid them and fail to gather the necessary feedback to correct this error. Conversely, if the costs are underestimated, the policy will visit those states and receive feedback, allowing for corrections. Therefore, smaller feedback buffer sizes can result in the learning of overly conservative policies.
Figure 12: Comparison of state level and trajectory level feedback elicitation.
Figure 13: Impact of the size of the feedback buffer on performance.
D.7 Need for the surrogate loss
Figure 14: Norm of the gradient when optimising the MLE loss v/s the proposed surrogate loss with trajectory level feedback.
In Section 4.2, we highlighted the challenges associated with minimizing Eq 4.2 particularly due to the term $\prod_{t=i}^{j}p^{safe}(s_{t},a_{t})$ which tends to collapse to 00 with longer segment lengths, resulting in unstable gradients. To investigate this further, we conducted a simple experiment in which we performed a few gradient steps to optimize Eq 4.2 and plotted the gradient norm. The results, shown in Figure 14 reveal that (a) directly optimizing Eq 4.2 produces gradients with a large norm that either quickly vanish or explode, rendering the learning process infeasible. In contrast, the surrogate objective in Eq 4.2 yields stable gradients allowing for convergence.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction are backed by theoretical and empirical results shown in Section 4.2 and 5 respectively.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of the proposed approach are discussed in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All the assumptions are discussed in Sections 3 and 4 and the complete proof can be found in the Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Apart from the Safety Gymansium environments that are publicly available, all other environments are described in detail so that they can be reproduced from scratch. The hyper parameters used to train the proposed approach are given in Appendix D.1 to ensure reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All the environments used are described in Section B. The code (with instructions on how to run) is publicly available: https://github.com/shshnkreddy/RLSF
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The hyperparameters used to run our propsed approach is present in Section D.1 in the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All results presented in the paper are averaged over multiple seeds and are presented with the standard error.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Details on the compute and time taken to produce the results is available in Section D.1 in the Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The presented research is compliant with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The social impacts of our work is discussed in Section 6.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The work presented in this paper has no risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The README file present in the code credits the authors, and the license and terms of use of has been properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The Safety Driver environments are described in Section B in the Appendix and instructions to run them are present in the README file included with the code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not conduct research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not conduct research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.